OpenStudy (anonymous):

express5√27 in the form n√3, where n is a positive interger

OpenStudy (across):

\[5\sqrt{27}=5\sqrt{3\cdot3^2}=15\sqrt{3}\]

OpenStudy (anonymous):

well done, across.

OpenStudy (anonymous):

n=15

OpenStudy (apoorvk):